OpenStudy (anonymous):

7^15/7^5 Simplify. A.1^3 B.7^10 C.7^20 D.7^3

OpenStudy (anonymous):

subtract \(15-5\) for your exponent

OpenStudy (anonymous):

B Because \[\frac{ 7^{15} }{ 7^{5} }=282475249\] and \[7^{10}=282475249\]. When dividing, the powers subtract, when multiplying, then add.

TheSmartOne (thesmartone):

\(\Huge \frac{x^y}{x^z}=x^{y-z}\)

OpenStudy (uri):